# Characterizing (multi-)semi-directed phylogenetic networks

### v1

@@ 1 Introduction
Phylogenetic networks are a generalization of evolutionary trees that are used to represent evolutionary histories of organisms such as plants and viruses that can evolve in a non-tree-like fashion (see e.g. huson2010phylogenetic ). In particular, they permit the representation of reticulate events, in which, for example, two species cross with one another or transfer genes. There are several classes of phylogenetic networks, but in this paper we will mainly focus on rooted phylogenetic networks (see e.g. kong2022classes for a recent overview), and some of their recent generalizations. Essentially, a rooted phylogenetic network is a directed acyclic graph, usually with a single source or root, whose leaf set corresponds to some collection of species or taxa. For example, in Figure 1, $D_{1}$ is a rooted phylogenetic network for the collection $\{x_{1},\dots,x_{8}\}$ of species. Lately, multi-rooted networks have also become of interest (see e.g. maxfield2025dissimilarity ; scholz2019osf ; soraggi2019general ), which only differ from rooted phylogenetic networks in that they may have multiple roots (see e.g. $D_{2}$ in Figure 1). In both types of networks, vertices with indegree greater than $1$ are of special interest because they represent reticulate events. Hence, such vertices are commonly called reticulations and their incoming arcs reticulation arcs.
@@
Recently, a new class of phylogenetic networks called semi-directed networks solis2016inferring has started to receive a lot of attention in the literature, both from a theoretical (e.g. banos2019identifying ; gross2018distinguishing ; gross2021distinguishing ; linz2023exploring ; xu2023identifiability ; englander2025identifiability ) and applied (e.g. allman2024nanuq+ ; holtgrefe2025squirrel ) point of view. Basically speaking, these are mixed graphs (i.e. graphs that can have a combination of undirected edges and directed arcs), which can be obtained by partly deorienting a rooted phylogenetic network, that is, by replacing all arcs which are not reticulation arcs with an edge, and then suppressing any resulting degree-2 vertex that arises from a root. We call such a partial deorientation, in which only reticulation arcs keep their direction and the root locations are lost, a semi-deorientation. For example, in Figure 1, $N_{1}$ is a semi-directed network since it is the semi-deorientation of, e.g., the rooted network $D_{1}$. In this paper, we will also consider multi-semi-directed networks, that is, mixed graphs which are the semi-deorientation of some multi-rooted network. As an example, in Figure 1, $N_{2}$ is multi-semi-directed since it is the semi-deorientation of $D_{2}$. In general, in case a (multi-)semi-directed network $N$ is a semi-deorientation of a (multi-)rooted network $D$, we shall call $D$ a rooting of $N$.
@@
Semi-directed networks have gained importance since there can be difficulties in locating the root in a phylogenetic network when inferring networks from biological data solis2016inferring ; gross2018distinguishing ; banos2019identifying ; gross2021distinguishing . However, their definition is somewhat problematic in that it is given implicitly rather than explicitly. For instance, the mixed graph $G$ in Figure 1 is neither semi-directed nor multi-semi-directed, but how can this be decided? Of course, one possibility would be to develop some algorithm to make this decision, but for certain applications it could also be useful to have characterizations for when a mixed graph is semi-directed or multi-semi-directed.
@@
In this paper, we shall provide some characterizations for semi-directed and multi-semi-directed networks. One of the main tools that we use to obtain our characterizations is the concept of a semi-directed cycle in a mixed graph, that is, a cycle of arcs and edges in the graph whose edges can be oriented so as to obtain a directed cycle. Indeed, the exclusion of such cycles is a condition in one of our main characterizations (see e.g. Theorem 2). As a corollary of our characterizations, in case a mixed graph is a (multi-)semi-directed network, we give a more general characterization to that given in maxfield2025dissimilarity for when a subset of vertices or subdivisions of edges or arcs can correspond to a choice of root(s) that leads to a rooting of the network.
@@
We shall also explore the consequences of our results for some special classes of rooted phylogenetic networks (see e.g. kong2022classes for a recent review of the different types of rooted networks). In particular, given a fixed class of rooted or multi-rooted phylogenetic networks, it is of interest to characterize when a (multi-)semi-directed network has a rooting contained in the given class. This could be either strongly (all rootings are in the class) or weakly (there exists a rooting that is in the class). For example, a rooted network is tree-based francis2015which if it has a rooted spanning tree with the same leaf set as the network. Thus, a semi-directed network is weakly tree-based if it has a rooting that is tree-based, and it is strongly tree-based if all its rootings are tree-based (see e.g. Figure 2).
[…]
@@
Various concepts have been used to characterize when a rooted phylogenetic network is contained within a certain class. For example, cherry picking sequences can be used for tree-child networks cardona2009comparison and orchards erdos2019class ; janssen2021cherry , omnians jetten2016nonbinary for tree-based networks francis2015which , and path partitions for forest-based networks huber2022forest . We shall generalize some of these concepts to (multi-)semi-directed networks and apply them to obtain characterizations for when such a network has a rooting that is contained within the classes of tree-child, orchard, tree-based or forest-based networks (or when all of its rootings are contained within these classes).
@@ 2.1 Mixed graphs
A mixed graph is an ordered tuple $G=(V,E,A)$ where $V$ is a nonempty set of vertices, $E$ is a set of undirected edges $\{u,v\}\subseteq V$, $u\neq v$, and $A$ is a set of directed arcs $(u,v)$ with $u,v\in V$, $u\neq v$, and such that for all arcs $(u,v)\in A$ we have that $\{u,v\}\notin E$ and $(v,u)\notin A$. Note that, by definition, parallel arcs, parallel edges or parallel edge/arc pairs are not allowed in mixed graphs. For an arc $(u,v)\in A$, we call $u$ the tail and $v$ the head. If $(u,v)\in A$, we call $u$ a parent of $v$ and $v$ a child of $u$. If there is an edge $\{u,v\}\in E$ or an arc $(u,v)\in A$, we call $u$ and $v$ adjacent or neighbours.
Suppose for the following that $G=(V,E,A)$ is a mixed graph. For $v\in V$, the indegree $d_{G}^{-}(v)$ is the number of arcs entering $v$, the outdegree $d_{G}^{+}(v)$ is the number of arcs leaving $v$, and $d_{G}^{e}(v)$ is the number of edges in $E$ incident to $v$. In addition, the degree $d_{G}(v)$ is the total number of edges and arcs incident to $v$. We will omit the subscript $G$ when the graph is clear from the context. We call $v$ a reticulation if $d^{-}(v)>1$, a leaf if $d(v)=d^{e}(v)=1$ or $d(v)=d^{-}(v)=1$ and a root if $d^{+}(v)=d(v)$. The set of leaves of $G$ is called the leaf set of $G$, and it is denoted by $L(G)$. We say that $G$ is binary if $d(v)\in\{1,3\}$ for all $v\in V$.
@@
A cycle in $G$ is a sequence of vertices $(v_{1},v_{2},\ldots,v_{p}=v_{1})$, $p\geq 4$, such that $v_{i}\neq v_{j}$ for $1\leq i<j<p$ and, for all $i\in\{1,\ldots,p-1\}$, either $(v_{i},v_{i+1})$ or $(v_{i+1},v_{i})$ is an arc in $A$ or $\{v_{i},v_{i+1}\}$ is an edge in $E$. Note that, since $p\geq 4$ and $v_{p}=v_{1}$, a cycle contains at least three distinct vertices. A cycle is called semi-directed if, for all $i\in\{1,\ldots,p-1\}$, either $(v_{i},v_{i+1})$ is an arc in $A$ or $\{v_{i},v_{i+1}\}$ is an edge in $E$. A reticulation $r$ of $N$ is a sink of a cycle in $N$ if $r=v_{i}$ and $(v_{i-1},v_{i}),(v_{i+1},v_{i})\in A$, for some $i\in\{1,\ldots,p-1\}$, with $v_{0}=v_{p-1}$. If $G$ is connected and does not contain a cycle then we call $G$ a tree. We refer to Figure 3 for examples that illustrate some of the definitions in this and the next subsection.
@@ 2.2 Multi-rooted and multi-semi-directed networks
Suppose $X$ is a finite set with at least two elements. A multi-rooted network (on $X$) is a mixed graph $(V,E,A)$ (with leaf set $X$), with $E=\emptyset$, no directed cycles, $d(v)\neq 2$ for all non-root vertices $v\in V$ and $d^{-}(v)\in\{0,d(v)-1\}$ for all $v\in V$. A $k$-rooted network is a multi-rooted network with precisely $k\geq 1$ roots. A $1$-rooted network is also called a rooted network. A rooted network without any reticulations is called a rooted phylogenetic tree.
@@
  • if $w$ has two incident arcs $(u,w),(w,v)$, replace them by a sinlge arc $(u,v)$;
@@
The semi-deorientation of a multi-rooted network $D$ is the result of replacing each arc $(u,v)$ of $D$ by an edge $\{u,v\}$ if $d_{D}^{-}(v)=1$ and afterwards suppressing any vertex $\rho$ with $d_{D}(\rho)=2$. Note that a root with two outgoing arcs in $D$ will still be present in the semi-deorientation. Also note that a semi-deorientation is not necessarily a mixed graph since suppressing roots may lead to parallel arcs. However, we will only consider mixed graphs in this paper and hence not consider cases with parallel arcs. A rooting [^1] Rootings were called “rooted partners” in maxfield2025dissimilarity ; linz2023exploring . of a mixed graph $G$ is a multi-rooted network $D$ such that $G$ is the semi-deorientation of $D$. Observe that $D$ can be obtained from $G$ by subdividing (zero or more) arcs and/or edges and replacing edges by arcs. Note that this includes the possibility that one or more leaves become roots. However, it is not possible to create any new reticulations. A $k$-semi-directed network (on $X$) is a mixed graph that is the semi-deorientation of some $k$-rooted network (with leaf set $X$). Note that the leaf set of a $k$-semi-directed network on $X$ may not be equal to $X$, because an outdegree-1 root may also be in $X$. A multi-semi-directed network is a $k$-semi-directed network for some $k\geq 1$. A semi-directed network is a $1$-semi-directed network. When drawing multi-rooted and multi-semi-directed networks, we will often omit the leaf labels when they are not relevant. We will reserve the letter $G$ for general mixed graphs, $D$ for (multi)-rooted networks and $N$ for (multi)-semi-directed networks.
@@ 3 Characterizations of multi-semi-directed networks
In this section we characterize when a mixed graph is a multi-semi-directed network or when it is in fact a semi-directed network.
@@
Figure 3: Left: A $2$-semi-directed network $N$ on $\{a,b,\ldots,e\}$ with set of reticulations $\{r_{1},r_{2},r_{3},r_{4}\}$. The sequence $(r_{1},v_{2},v_{5},r_{2},v_{3},v_{4},r_{3},r_{4},g)$ is a $\wedge$-path of $N$ and the sequence $C=(r_{2},v_{3},v_{4},r_{3},r_{4},v_{7},r_{2})$ is a cycle of $N$. The vertices $r_{2}$ and $r_{4}$ are sinks of $C$ whereas $r_{3}$ is not. The path $(a,v_{1},v_{3},v_{4},v_{7})$ is an example of an edge-path. The leaves $a$ and $b$ form a cherry, leaves $d$ and $g$ are reticulation leaves, while $r_{1}$ and $r_{4}$ are leaf reticulations. The vertex sets of the source components are $\{a,b,v_{1},v_{3},v_{4},v_{7}\}$ and $\{e\}$ while the sink components have vertex sets $\{d,r_{1}\}$ and $\{g,r_{4}\}$. Right: A rooting $D$ of $N$ in the form of a $2$-rooted network with roots $\rho_{1}$ and $\rho_{2}$.
@@
Let $C=(v_{1},v_{2},\ldots,v_{p}=v_{1})$, $p\geq 4$. If $C$ is not semi-directed then there exist $i,j\in\{1,2,\ldots,p-1\}$ distinct such that $(v_{i},v_{i+1})$ and $(v_{j+1},v_{j})$ are arcs on $C$. Without loss of generality, $i<j$ $(\mathrm{mod}\,\,\,\,p-1)$. Assume that $j$ and $i$ are such that $j-i$ is minimized. If $j-i=1$ then $C$ contains the sink $v_{j}=v_{i+1}$. Otherwise, $j-i>1$ and, so, $v_{i+1}$ and $v_{j}$ are both reticulations of $G$ and $C$ contains a non-trivial edge-path between them. ∎

### v2

# Characterizing semi-directed phylogenetic networks and their multi-rootable variants

@@ 1 Introduction
Phylogenetic networks are a generalization of evolutionary trees that are used to represent evolutionary histories of organisms such as plants and viruses that can evolve in a non-tree-like fashion (see e.g. huson2010phylogenetic ). In particular, they permit the representation of reticulate events, in which, for example, two species cross with one another or transfer genes. There are several classes of phylogenetic networks, but in this paper we will mainly focus on rooted phylogenetic networks (see e.g. kong2022classes for a recent overview), and some of their recent generalizations. Essentially, a rooted phylogenetic network is a directed acyclic graph, usually with a single source or root, whose leaf set corresponds to some collection of species or taxa. For example, in Figure 1, $D_{1}$ is a rooted phylogenetic network for the collection $\{x_{1},\dots,x_{8}\}$ of species. Lately, multi-rooted networks have also become of interest, which only differ from rooted phylogenetic networks in that they may have multiple roots (see e.g. $D_{2}$ in Figure 1). Such networks can be used to model ancestral relationships between populations soraggi2019general and for representing the evolutionary history of distantly related groups of species that can still exchange genes huber2022forest ; scholz2019osf . In both types of networks, vertices with indegree greater than $1$ are of special interest because they represent reticulate events. Hence, such vertices are commonly called reticulations and their incoming arcs reticulation arcs.
@@
Recently, a new class of phylogenetic networks called semi-directed networks solis2016inferring has started to receive a lot of attention in the literature, both from a theoretical (e.g. banos2019identifying ; gross2018distinguishing ; gross2021distinguishing ; linz2023exploring ; xu2023identifiability ; englander2025identifiability ) and applied (e.g. allman2024nanuq+ ; holtgrefe2025squirrel ) point of view. Basically speaking, these are mixed graphs (i.e. graphs that can have a combination of undirected edges and directed arcs), which can be obtained by partly deorienting a rooted phylogenetic network, that is, by replacing all arcs which are not reticulation arcs with an edge, and then suppressing any resulting degree-2 vertex that arises from a root. We call such a partial deorientation, in which only reticulation arcs keep their direction and the root locations are lost, a semi-deorientation. For example, in Figure 1, $N_{1}$ is a semi-directed network since it is the semi-deorientation of, e.g., the rooted network $D_{1}$. In this paper, we will also consider multi-semi-directed networks, that is, mixed graphs which are the semi-deorientation of some multi-rooted network. Note that these networks were recently considered in maxfield2025dissimilarity in the context of defining a dissimilarity measure between semi-directed networks. As an example, in Figure 1, $N_{2}$ is multi-semi-directed since it is the semi-deorientation of $D_{2}$. In general, in case a (multi-)semi-directed network $N$ is a semi-deorientation of a (multi-)rooted network $D$, we shall call $D$ a rooting of $N$.
@@
Locating the root in a phylogenetic network inferred from biological data is often problematic. In particular, without asymmetrical models of character change, root placement cannot be determined from the data alone and must rely on external assumptions kinene2016rooting . Due to more favorable identifiability results that circumvent the need for root placement solis2016inferring ; gross2018distinguishing ; banos2019identifying ; gross2021distinguishing , semi-directed networks have gained importance instead. However, their definition is somewhat problematic, in that it is given implicitly rather than explicitly. For instance, the mixed graph $G$ in Figure 1 is neither semi-directed nor multi-semi-directed, but how can this be decided? Of course, one possibility would be to develop some algorithm to make this decision, but for certain applications it could also be useful to have characterizations for when a mixed graph is semi-directed or multi-semi-directed.
@@
In this paper, we shall provide some characterizations for semi-directed and multi-semi-directed networks. These results reveal combinatorial features of (multi-)semi-directed networks that help support further results in this paper and may also be helpful for future studies. One of the main tools that we use to obtain our characterizations is the concept of a semi-directed cycle in a mixed graph, that is, a cycle of arcs and edges in the graph whose edges can be oriented so as to obtain a directed cycle. Indeed, the exclusion of such cycles is a condition in one of our main characterizations (see e.g. Theorem 2). As a corollary of our characterizations, in case a mixed graph is a (multi-)semi-directed network, we give a more general characterization to that given in maxfield2025dissimilarity for when a subset of vertices or subdivisions of edges or arcs can correspond to a choice of root(s) that leads to a rooting of the network.
@@
We shall also explore the consequences of our results for some special classes of rooted phylogenetic networks (see e.g. kong2022classes for a recent review of the different types of rooted networks). In particular, given a fixed class of rooted or multi-rooted phylogenetic networks, it is of interest to characterize when a (multi-)semi-directed network has a rooting contained in the given class. This could be either strongly (all rootings are in the class) or weakly (there exists a rooting that is in the class). For example, a rooted phylogenetic network is tree-based francis2015which if it has a rooted spanning tree with the same leaf set as the network. Thus, a semi-directed network is weakly tree-based if it has a rooting that is tree-based, and it is strongly tree-based if all its rootings are tree-based (see e.g. Figure 2).
@@
Various concepts have been used to characterize when a rooted phylogenetic network is contained within a certain class. For example, the classes of rooted tree-child networks cardona2009comparison and rooted orchard networks erdos2019class ; janssen2021cherry can both be characterized using cherry picking sequences erdos2019class ; janssen2021cherry , eventually leading to practical software for reconciling phylogenetic trees into networks bernardini2023constructing ; bernardini2024inferring . On the other hand, rooted tree-based networks francis2015which have been characterized using both omnians jetten2016nonbinary and path partitions francis2018new , with the latter concept also providing a characterization for rooted forest-based networks huber2022forest . We shall generalize some of these concepts to (multi-)semi-directed networks and apply them to obtain characterizations for when such a network has a rooting that is contained within the classes of tree-child, orchard, tree-based or forest-based networks (or when all of its rootings are contained within these classes).
@@ 2.1 Mixed graphs
A mixed graph is an ordered tuple $G=(V,E,A)$ where $V$ is a nonempty set of vertices, $E$ is a set of undirected edges $\{u,v\}\subseteq V$, $u\neq v$, and $A$ is a set of directed arcs $(u,v)$ with $\in V, uvu\neq v$, and such that for all arcs $(u,v)\in A$ we have that $\{u,v\}\notin E$ and $(v,u)\notin A$. Note that, by definition, parallel arcs, parallel edges or parallel edge/arc pairs are not allowed in mixed graphs. For an arc $(u,v)\in A$, we call $u$ the tail and $v$ the head. If $(u,v)\in A$, we call $u$ a parent of $v$ and $v$ a child of $u$. If there is an edge $\{u,v\}\in E$ or an arc $(u,v)\in A$, we call $u$ and $v$ adjacent or neighbours.
Suppose for the following that $G=(V,E,A)$ is a mixed graph. For $v\in V$, the indegree $d_{G}^{-}(v)$ is the number of arcs entering $v$, the outdegree $d_{G}^{+}(v)$ is the number of arcs leaving $v$, and $d_{G}^{e}(v)$ is the number of edges in $E$ incident to $v$. In addition, the degree $d_{G}(v)$ is the total number of edges and arcs incident to $v$. We will omit the subscript $G$ when the graph is clear from the context. We call $v$ a reticulation if $d^{-}(v)>1$, a leaf if $d(v)=d^{e}(v)=1$ or $d(v)=d^{-}(v)=1$ and a root if $d^{+}(v)=d(v)$. The set of leaves of $G$ is called the leaf set of $G$, and it is denoted by $L(G)$. We say that $G$ is binary if $d(v)\in\{1,2\}$ for each root $v\in V$ and $d(v)\in\{1,3\}$ for each non-root $v\in V$.
@@
A cycle in $G$ is a sequence of vertices $(v_{1},v_{2},\ldots,v_{p}=v_{1})$, $p\geq 4$, such that $v_{i}\neq v_{j}$ for $1\leq i<j<p$ and, for all $i\in\{1,\ldots,p-1\}$, either $(v_{i},v_{i+1})$ or $(v_{i+1},v_{i})$ is an arc in $A$ or $\{v_{i},v_{i+1}\}$ is an edge in $E$. Note that, since $p\geq 4$ and $v_{p}=v_{1}$, a cycle contains at least three distinct vertices. A cycle is called semi-directed if, for all $i\in\{1,\ldots,p-1\}$, either $(v_{i},v_{i+1})$ is an arc in $A$ or $\{v_{i},v_{i+1}\}$ is an edge in $E$. A reticulation $r$ of $N$ is a sink of a cycle in $N$ if $r=v_{i}$ and $(v_{i-1},v_{i}),(v_{i+1},v_{i})\in A$, for some $i\in\{1,\ldots,p-1\}$, with $v_{0}=v_{p-1}$. If $G$ is connected and does not contain a cycle then we call $G$ a tree. If additionally $E=\emptyset$, i.e., $G$ is fully directed, and $G$ has a single root, we call $G$ a rooted tree. We refer to Figure 3 for examples that illustrate some of the definitions in this and the next subsection.
@@ 2.2 Multi-rooted and multi-semi-directed networks
Suppose $X$ is a finite set with at least two elements. A multi-rooted network (on $X$) is a mixed graph $(V,E,A)$ (with leaf set $X$), with $E=\emptyset$, no directed cycles, $d(v)\neq 2$ for all non-root vertices $v\in V$ and $d^{-}(v)\in\{0,1,d(v)-1\}$ for all $v\in V$. A $k$-rooted network is a multi-rooted network with precisely $k\geq 1$ roots. A $1$-rooted network is also called a rooted network. A rooted network without any reticulations is called a rooted phylogenetic tree.
@@
  • if $w$ has two incident arcs $(u,w),(w,v)$, replace them by a single arc $(u,v)$;
@@
The semi-deorientation of a multi-rooted network $D$ is the result of replacing each arc $(u,v)$ of $D$ by an edge $\{u,v\}$ if $d_{D}^{-}(v)=1$ and afterwards suppressing any vertex $\rho$ with $d_{D}(\rho)=2$. Note that a root with two outgoing arcs in $D$ will still be present in the semi-deorientation. Also note that a semi-deorientation is not necessarily a mixed graph since suppressing roots may lead to parallel arcs. However, we will only consider mixed graphs in this paper and hence not consider cases with parallel arcs. A rooting [^2] Rootings were called “rooted partners” in maxfield2025dissimilarity ; linz2023exploring . of a mixed graph $G$ is a multi-rooted network $D$ such that $G$ is the semi-deorientation of $D$. Observe that $D$ can be obtained from $G$ by subdividing (zero or more) arcs and/or edges and replacing edges by arcs. Note that the subdivision vertices necessarily become roots in $D$ and that vertices of $G$ may also become roots in $D$. This also includes the possibility that a root of $G$ that is contained in $X$ becomes a leaf in $D$ (see e.g. the vertex $e$ in Figure 3) or that a leaf of $G$ not contained in $X$ becomes an outdegree-1 root in DD [^3] From a biological perspective, turning a leaf in a (multi-)semi-directed network into a root of a (multi)-rooted network may seem problematic. However, since this only applies to leaves of a (multi-)semi-directed network not in $X$, which do not correspond to taxa, this does not pose a problem. A leaf of a (multi-)semi-directed network contained in $X$ cannot be turned into a root of a (multi)-rooted network. Instead, its incident edge can be subdivided by a new vertex that becomes a root. . However, it is not possible to create any new reticulations. A $k$-semi-directed network (on $X$) is a mixed graph that is the semi-deorientation of some $k$-rooted network (with leaf set $X$). Note that the leaf set of a $k$-semi-directed network on $X$ may not be equal to $X$, because an outdegree-1 root may also be in $X$ (see, for example, the vertex $e$ in Figure 3 again). A multi-semi-directed network is a $k$-semi-directed network for some $k\geq 1$. A semi-directed network is a $1$-semi-directed network. When drawing multi-rooted and multi-semi-directed networks, we will often omit the leaf labels when they are not relevant. We will reserve the letter $G$ for general mixed graphs, $D$ for (multi)-rooted networks and $N$ for (multi)-semi-directed networks.
@@ 3 Characterizations of multi-semi-directed networks
In this section we characterize when a mixed graph is a multi-semi-directed network or when it is in fact a semi-directed network. These characterizations naturally lead to algorithms for checking (multi-)semi-directedness. We sketch one efficient approach at the end of Section 4, which outlines a method based on rootings of (multi-)semi-directed networks.
@@
Figure 3: Left: A $2$-semi-directed network $N$ on $\{a,b,\ldots,e\}$ with set of reticulations $\{r_{1},r_{2},r_{3},r_{4}\}$. The sequence $(r_{1},v_{2},v_{5},r_{2},v_{3},v_{4},r_{3},r_{4},g)$ is a $\wedge$-path of $N$ and the sequence $C=(r_{2},v_{3},v_{4},r_{3},r_{4},v_{7},r_{2})$ is a cycle of $N$. The vertices $r_{2}$ and $r_{4}$ are sinks of $C$ whereas $r_{3}$ is not. The path $(a,v_{1},v_{3},v_{4},v_{7})$ is an example of an edge-path. The leaves $a$ and $b$ form a cherry, leaves $d$ and $g$ are reticulation leaves, while $r_{1}$ and $r_{4}$ are leaf reticulations. The vertex sets of the source components are $\{a,b,v_{1},v_{3},v_{4},v_{7}\}$ and $\{e\}$ while the sink components have vertex sets $\{d,r_{1}\}$ and $\{g,r_{4}\}$. Right: A rooting $D$ of $N$ in the form of a 22-rooted network with roots $\rho_{1}$ and $\rho_{2}$.
@@
Let $C=(v_{1},v_{2},\ldots,v_{p}=v_{1})$, $p\geq 4$. If $C$ is not semi-directed then there exist $i,j\in\{1,2,\ldots,p-1\}$ distinct such that $(v_{i},v_{i+1})$ and $(v_{j+1},v_{j})$ are arcs on $C$. Without loss of generality, $i<j$ $(\mathrm{mod}\,\,\,\,p-1)$. Assume that $j$ and $i$ are such that jij-i is minimized. If ji=1j-i=1 then $C$ contains the sink $v_{j}=v_{i+1}$. Otherwise, ji>1j-i>1 and, so, $v_{i+1}$ and $v_{j}$ are both reticulations of $G$ and $C$ contains a non-trivial edge-path between them. ∎